LUCIFER Technical Observing, prep for Partner Science: UT 2010 January 04

Observer: DThompson, MPedani (LBTO)
Telescope Operator: WWack
Instrument Support: DOfficer, WSeifert (on call)
Telescope Support: JUrban
SW support: TLeibold (on call)

Summary

  • Completed Garnavich SDT H band imaging on SN2008Q
  • Implemented LUCIFERSX.20100103.ptmod as default
  • Successfully guided on stars to 18.3 mag in 0.5 arcsec seeing
  • Successfully collimated on 18 mag star with 120s exposure
    • Need to work on centroiding and thresholding algorithms in wfs.
  • Collected a better set of field aberration measurements in good seeing

Preparations

- Analysis of field aberration data pending, but Doug's quick look suggests no large problems.

Details

0:00 Start twilight flats:
01:03 Correct initial collimation/pointing with IE/CA

01:19 Garnavich field H band luci.20100104.0034 - 0083

03:00 Slew to Stone-C Field to test Guide/WFS down to to R~18.3 mag (DSS2 mag)

Stone-C GS=05 R=14.36 luci.20100104.0084 Ks 10s x 6
Stone-C GS=18 R=14.36 luci.20100104.0085 Ks 10s x 6
Stone-C GS=21 R=16.19 luci.20100104.0086 Ks 10s x 6
Stone-C GS=19 R=17.26 luci.20100104.0087/88 Ks 10s x 6. We are not WFS because of a threshold in the WFS algorithm; guide is OK
Stone-C GS=06 R=17.71 luci.20100104.0089/90 Ks 10s x 6. We are not WFS but we collimated on GS=21 first; guide is OK
Stone-C GS=28 R2=17.99 luci.20100104.0091/92 Ks 10s x 6. We are not WFS but we collimated on GS=21 first; guide is OK
Stone-C GS=29 R2=18.30 luci.20100104.0093/94 Ks 10s x 6. We are not WFS but we collimated on GS=21 first; guide is OK (seeing ~0.6")

04:38 Stone-C Field LUCI Rot center Measurement

Stone-C GS=17 PA=-70 luci.20100104.0095 Ks 2s x 5
Stone-C GS=17 PA=-35 luci.20100104.0096 Ks 2s x 5
Stone-C GS=17 PA=0 luci.20100104.0097 Ks 2s x 5
Stone-C GS=17 PA=+35 luci.20100104.0098 Ks 2s x 5
Stone-C GS=17 PA=+70 luci.20100104.0099 Ks 2s x 5

05:00 Check that both guide and WFS work on faint stars with 120s WFS exposures
Changed in AIP_L.cfg file:
  • WFScam_exposureTime 120000 (from 30000)
  • WFScam_maxExposure 120000 (from 30000)
  • countThreshold 50 (from 100)
and sent preset to the 18th mag guide star in StoneCagw field. WFS images 219-222 are still 30s.

Now Exp. Time is 120s and we proceed.
Stone-C GS=21 R=16.19 luci.20100104.0100 Ks 10s x 6
Stone-C GS=28 R2=17.99 luci.20100104.0101-105 Ks 10s x 6. We are WFS now!; guide is OK; seeing is ~0.5"

05:26 Reset AIP_L.cfg exposure times to previous values (left countThreshold at 50), re-authorized LUCIFER

05:39 Slew to BS9107 to take a better set of field aberration data
script Field_BS9107.cl copied to: /OldRepository/AGW-Data/100104_BS9107
Note that we forgot to switch to "primary" mode - data taken in "split" mode. I hope that is OK.

GS# wfsc#
0 03-04
82 05-07
0 08-09
35 10-12
0 13-14
19 15-17
0 18-19
64 20-22
0 23-24
96 25-27
0 28-29
40 30-32
0 33-34
35 35-37
0 38-39
48 40-42
0 43-44
8 45-47
0 48-49
68 50-52
0 53-54
96 55-57
0 58-59
70* 60-62
0 63-64
39 65-67
0 68-69
97 70-72
0 73-74
70 75-77
0 78-79
82 80-82
0 83-84
35 85-87
0 88-89

Rotator went into velocity mode on preset to GS70. Preset failed. Script continued, but frames 60-62 have no star on the WFS. Wayne recovered by putting the rotator in hold and then back to track mode so the script was not interrupted. The next preset to GS0 worked fine. Seeing is still ~0.4 arcsec on the guider. Note that the guider images frequently looked elongated while these data were being taken. Supplementary data taken with script field_BS9107_supp.cl (retake GS70 plus repeat the first 4 in the original script).

07:47 MCSPU hung in preset to GS19? 16 good points taken in good seeing. Called Torsten - no answer. Called Norm - he said to call Tom Sargent. Tom said he would have a look. Recovery took time, so we are done with the field aberration measurements. See IT# Jet apparently having trouble talking with the DSP board. Back up 08:27, lost 40 minutes

08:28:57.467 2010 Preset alert for left LUCIFER mode TRACK: Invalid buffer: RPCERROR. Sending preset hung on IIF at "Sending target (PCS)". Stopped and restarted PCS.

Then on next preset PCS reports MCSPU error returned - ERROR rotator is not movable (Not 'READY' for some reason).

Check pointing on BS9107 before resetting encoders...only IE needed offset of +8 arcsec.

08:33 reset encoders

08:53 Slew to Bik SDT Target NS2_3 to Practice spectroscopic acquisitions
Acquisition frames: 0107(slit) and 0108 (image)
Spectra: luci.20100104.0109-0112. Seeing ~0.6-0.8 and variable on guider. First spectrum is 0.63 arcsec FWHM in the spatial direction.
Arcs: 0113-0124 Flats: 0125-0128

09:48 Slew to Bik SDT Target STD255_Disk
Spectra: luci.20100104.0133-0134; star well out of focus. we were not collimated.

10:09 Slew to Bik SDT Target S255_Disk

At 10:12 the Instrument Control GUI remains Yellow even if the setup was successful; we re-initialize; Successful.

During the acquisition the image: luci.20100104.0138 comes out of focus despite we were collimating; some jump ? Spectra: luci.20100104.0141-144

10:52 Slew to Bik SDT Target STD255_Disk
Spectra: luci.20100104.0150-0151

11:10 Slew to McLinden SDT Target one_40844
Oops. It's only 4 deg from Moon. We skip it !

11:19 Slew to Assef SDT Target SDSS1138
Problem with Acquisition Setup; The script tried to put the NB6 mask (0.5") in the beam but it has been moved to NB26 now! Called Walter. Initialize and fixed.

11:44 we get a AGW timeout after sending a 10" +x offset Faint Target! We edited the script to acquisition offset +5 +5 then re-acquired the target at 0,0 to subctract.

Spectra: luci.20100104.0158-0161 The 0.5" slit has gaps! At least the 0161 fell into it. We modify the script with the following offset at Y: +30 -5 -5 -5 and re-launch the script.

Spectra: luci.20100104.0162-0165 Flats: 0167

12:44 check pointing ACT0205 but we get an error while clearing the beam for imaging 2010.01.04 12:44:25 s | Error in sequence MOS - FPU to turnout. WARNING holding current still active at [card number: 1, motor number: 2, type: "mask handling unit head rotation", description: "rotator"]!
Call Walter...problem with a switch that did not get tripped when it was supposed to. Details will follow from Walter and Peter. Recovered at 13:30 (lost 45 minutes). It is late enough that we cannot do anything else, so we are putting in the blind mask/filters to do some darks.

14:01 Start taking Darks

Files: luci.20100104.0168-0212

Observing Plan

* Correct initial collimation/pointing with IE/CA
* Restest updatePointingReference functionality
* Collect astrometric field data to check telescope scale at colder temps (use science data for this)

* Take a second set of pointing data to combine in TPOINT
  • Reset mount encoders (now takes only 2 minutes) @AZ=120, EL=90
  • Have TO start the pointing log!
  • Check first position in AZ < 0 degrees wrap
  • Use ARC method to measure AGw/Rotator center
  • 28 points are sufficient, more is always better
  • Can use NB filter in LUCIFER, the pointing stars are ~7.5 mag
  • needs stable temps!
  • May need to do several short PMs as temps are changing rapidly

* Re-Take Field aberration measurements if necessary (to confirm new numbers?)

* Demonstrate guiding to 18th mag (Tel=GUIDE)
* WFS on 18th mag. Will likely need to set the wfs max exposure time to 120s
* Quantify the guide/wfs limits with mag and seeing vs. LUCIFER DIQ

-- DavidThompson - 03 Jan 2010
Topic attachments
I Attachment Action Size Date Who Comment
daq.20100104.jpgjpg daq.20100104.jpg manage 609 K 07 Jan 2010 - 00:54 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.sx.thermal.20100104.jpgjpg pmcc.sx.thermal.20100104.jpg manage 429 K 07 Jan 2010 - 00:54 JohnHill Plot of SX Primary Mirror Temperatures
zern.sx.20100104.gifgif zern.sx.20100104.gif manage 21 K 12 Jan 2010 - 21:45 JohnHill Plot of Z11 and Z22 corrections
Topic revision: r15 - 12 Jan 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback